OpenStudy (anonymous):

please prove this by induction for all non-negative integers n. 1/2 + 2/(2^2) +3/(2^3) +...+n/(2^n) = 2- (n+2/(2^n))

OpenStudy (campbell_st):

start by proving it true for n = 1 1st term is 1/2 now substitute n = 1 into the right hand side of the equation.

OpenStudy (campbell_st):

and show the sum of 1 term is 1/2

OpenStudy (campbell_st):

that is the 1st step to induction

OpenStudy (campbell_st):

the next step is to assume its true for a value n = k so the sum of k terms is \[2 - (\frac{k + 2}{2^k})\]

OpenStudy (anonymous):

thanks. what about the last step n=k+1 ?

OpenStudy (campbell_st):

yes so find the k + 1 term and add it to the sum of k terms. then show it is equal to the sum of k + 1 terms.

OpenStudy (campbell_st):

so you need to show k +1 term is \[\frac{k + 1}{2^{k + 1}}\] then show \[S_{k} + t_{k + 1} = S_{k + 1}\]

OpenStudy (campbell_st):

oops should read \[2 = \frac{k + 2}{2^k} + \frac{k + 1}{2^{k + 1}} = 2 - \frac{(k + 1) + 2}{2^{k + 1}}\]

OpenStudy (campbell_st):

dang... it should read 2 - ... not 2 =

OpenStudy (anonymous):

thanks a lot. really helping

OpenStudy (campbell_st):

its reasonably easy from here... just be careful with the signs of terms... and use the fact that \[2^k = \frac{2}{2^{k + 1}}\]